# MP5: A Multi-modal Open-ended Embodied System in Minecraft via Active Perception

### v3

@@ 4.3 Ablation Study
We conduct ablation studies to evaluate the effectiveness of various modules. The experimental setup and the associated success rates are in Sec. 4.1. More detailed ablation studies are listed in Sup.C.3. The following paragraphs present the analyses derived from our ablation studies.
@@
Robustness is essential in open-world settings. To enhance the robustness evaluation of our system, we introduce a “Random Drop” setting. In this setting, we randomly discard one complete sub-objective from the inventory at the start of each new sub-objective, which deliberately induces execution errors for the agent. Comparing the second and third lines in Tab. 7, we observe the critical role of the Patroller in recognizing feedback errors. The Patroller’s ability to integrate current environmental information with error information is essential for enabling the planner to re-plan. The significance of this robustness is evident when examining the success rates. Without the Patroller’s robustness, the agent’s success rate on the Wooden level [Uncaptioned image] plummets from $76.67\%$ to $7.33\%$, while success rates on the Iron [Uncaptioned image], and Diamond [Uncaptioned image] levels drop to $0.00\%$. Details regarding the “Random Drop” setting can be found in Sup.C.3.
@@ 5 Conclusion
  • Sec. A: Percipient, Memory, Observation Space and Action Space.
  • Sec. B: Environment Setting.
  • Sec. C: Task Details, Sucess Rates of All Tasks and Ablation Study.
  • Sec. D: Different Strategy of Active Perception.
  • Sec. E: Applications Demonstration of MP5.
  • Sec. F: Interactions in MP5.
Appendix A Implementation Details
@@
A.1 Percipient
A.1.1 Data Collection
@@
A.1.2 MineLLM training details
MineLLM combines the image visual encoder from MineCLIP [9] and the large language models from Vicuna-13B-v1.5 [5]. Images are processed by the frozen vision encoder, whose features are projected by a two-layer MLP named Alignment Net to the same feature space as the text embeddings of the applied LLM. Instructions are tokenized by SentencePiece tokenizer [15], and then the vision and text tokens are concatenated to feed into the LLM model. To better align the feature space of visual image encoder from MineCLIP [9] and large language model from Vicuna [5], we collect $500K$ image-text instruction-following data on the MineDojo [9] following the method detailed in Section A.1.1, for the purpose of training MineLLM. Each training instance consists of an image $\mathcal{I}$ and a multi-turn conversation data $(\boldsymbol{x}_{1},\boldsymbol{y}_{1},\ldots,\boldsymbol{x}_{n},\boldsymbol{y}_{n})$, where $\boldsymbol{x}_{i}$ and $\boldsymbol{y}_{i}$ are the human’s instruction and the system’s response at the $i$-th turn. To train MineLLM efficiently, we add LoRA [13] parameters to all projection layers of the self-attention layers in the LLM. Only the parameters of the Alignment Net and the LoRA [13] module are optimized during training. Multimodal tokens are decoded by the LLM model and the corresponding LoRA [13] parameters.
@@ A.4 Action Space
Appendix B Environment Setting
@@ 3-4 Find a pig [Uncaptioned image] during the nighttime [Uncaptioned image] in a rainy day
To better evaluate Context-Dependent Tasks and Process-Dependent Tasks, which are defined in detail in Section C.1, we select different environment settings in MineDojo [9]. For Context-Dependent Tasks, we uniformly adopt the environment in MineDojo [9] with the creative “task_id” of “0”. For Process-Dependent Tasks, we uniformly adopt the environment with the “task_id” of “harvest”, “target_names” as “diamond”, and “spawn_rate” as “1”. This is why obtaining redstone is more difficult than obtaining diamond, as described in Section C.2.2.
Appendix C Task Details and Experiment Results
@@
C.1 Task Details
C.1.1 Context-Dependent Tasks
@@
C.1.2 Process-Dependent Tasks
@@
C.2 Success Rates of All Tasks
C.2.1 Context-Dependent Tasks
@@
[…]
@@
C.2.2 Process-Dependent Tasks
@@
C.3 Ablation Study
C.3.1 Context-Dependent Tasks
We conduct ablation studies on the multi-modal large language model (MLLM) part within Context-Dependent Tasks in 14, comparing the performance outcomes of different MLLMs and different pre-trained visual encoders in the percipient.
C.3.2 Process-Dependent Tasks
@@
Appendix D Different Strategy of Active Perception
@@
Appendix E Applications
E.1 Obtain Diamond Pickaxe
@@ E.2 Discovery
E.3 Open-Ended Tasks

### v4

@@ 4.3 Ablation Study
We conduct ablation studies to evaluate the effectiveness of various modules. The experimental setup and the associated success rates are in Sec. 4.1. More detailed ablation studies are listed in Sup.D.3. The following paragraphs present the analyses derived from our ablation studies.
@@
Robustness is essential in open-world settings. To enhance the robustness evaluation of our system, we introduce a “Random Drop” setting. In this setting, we randomly discard one complete sub-objective from the inventory at the start of each new sub-objective, which deliberately induces execution errors for the agent. Comparing the second and third lines in Tab. 7, we observe the critical role of the Patroller in recognizing feedback errors. The Patroller’s ability to integrate current environmental information with error information is essential for enabling the planner to re-plan. The significance of this robustness is evident when examining the success rates. Without the Patroller’s robustness, the agent’s success rate on the Wooden level [Uncaptioned image] plummets from $76.67\%$ to $7.33\%$, while success rates on the Iron [Uncaptioned image], and Diamond [Uncaptioned image] levels drop to $0.00\%$. Details regarding the “Random Drop” setting can be found in Sup.D.3.
@@ 5 Conclusion
Acknowledgement. This work was supported by the National Key R&D Program of China (2021YFB1714300), the National Natural Science Foundation of China (62106154, 62132001), the Natural Science Foundation of Guangdong Province, China (2022A1515011524).
@@
  • Sec. A: Discussion of MP5.
  • Sec. B: Percipient, Memory, Observation Space and Action Space.
  • Sec. C: Environment Setting.
  • Sec. D: Task Details, Sucess Rates of All Tasks and Ablation Study.
  • Sec. E: Different Strategy of Active Perception.
  • Sec. F: Applications Demonstration of MP5.
  • Sec. G: Interactions in MP5.
Table 8: Explicit comparisons of setups and consequences.
Method Observation Space Action Space Instruct Situation- Situation- Tasks Agents Can Perform
Info Not Action Action Primitive Following aware aware Process Context Process & Context
Type Omniscient Type Num Library Size Plan Execution (Long-Horizon Tasks) (High Env. Info Tasks) (Combination of 2 Tasks)
DreamerV3 RGB Original MineRL 25 0
DEPS RGB Compound (MineDojo) 42 0
GITM Text Manual (MineDojo) 9 9 + corner cases
Voyager Text JavaScript APIs N/A All
MP5(ours) RGB Compound (MineDojo) 10 4
Appendix A Discussion of MP5
As shown in Fig. 6, existing methods usually follow the paradigm that an agent should be with the ability of planning (e.g., GITM & Voyager), embodiment (e.g., DreamerV3), or both of them (e.g., DEPS). To be different, MP5 introduces a new paradigm that these components (i.e., embodiment, planning, and execution as well), should be enhanced with the awareness of the situation (i.e., rich contextual and procedural information w.r.t. the task), which is more human-like and has the potential to solve more difficult open-end context- and process-dependent tasks (as evaluated in Sec. 4.2.3). (2) Technically, MP5 comprises five interactive modules to meet the requirement of the new paradigm, with an MLLM-based active perception scheme to fulfil the situation awareness. To our best knowledge, MP5 is the first embodied system in Minecraft that is capable of situation-aware planning and action execution. (3) Only our method can solve Context-Dependent Tasks (in Tab. 8) by leveraging the situation-aware planning and execution. We constructed a detailed benchmark (in Sec. 4) to explore how agents complete tasks required complex reasoning and constrained by extensive environmental information.
We have provided a comparison of setups and their consequences in Tab. 8. It tells that the proposed MP5 applies ego-centric RGB images, just utilizes a restrained set of human-defined primitives, but can solve the most challenging context-dependent and process-dependent tasks.
Figure 6: Paradigm innovation.
Appendix B Implementation Details
@@
B.1 Percipient
B.1.1 Data Collection
@@
B.1.2 MineLLM training details
MineLLM combines the image visual encoder from MineCLIP [9] and the large language models from Vicuna-13B-v1.5 [5]. Images are processed by the frozen vision encoder, whose features are projected by a two-layer MLP named Alignment Net to the same feature space as the text embeddings of the applied LLM. Instructions are tokenized by SentencePiece tokenizer [16], and then the vision and text tokens are concatenated to feed into the LLM model. To better align the feature space of visual image encoder from MineCLIP [9] and large language model from Vicuna [5], we collect $500K$ image-text instruction-following data on the MineDojo [9] following the method detailed in Section B.1.1, for the purpose of training MineLLM. Each training instance consists of an image $\mathcal{I}$ and a multi-turn conversation data $(\boldsymbol{x}_{1},\boldsymbol{y}_{1},\ldots,\boldsymbol{x}_{n},\boldsymbol{y}_{n})$, where $\boldsymbol{x}_{i}$ and $\boldsymbol{y}_{i}$ are the human’s instruction and the system’s response at the $i$-th turn. To train MineLLM efficiently, we add LoRA [13] parameters to all projection layers of the self-attention layers in the LLM. Only the parameters of the Alignment Net and the LoRA [13] module are optimized during training. Multimodal tokens are decoded by the LLM model and the corresponding LoRA [13] parameters.
@@ B.4 Action Space
Appendix C Environment Setting
@@ 3-4 Find a pig [Uncaptioned image] during the nighttime [Uncaptioned image] in a rainy day
To better evaluate Context-Dependent Tasks and Process-Dependent Tasks, which are defined in detail in Section D.1, we select different environment settings in MineDojo [9]. For Context-Dependent Tasks, we uniformly adopt the environment in MineDojo [9] with the creative “task_id” of “0”. For Process-Dependent Tasks, we uniformly adopt the environment with the “task_id” of “harvest”, “target_names” as “diamond”, and “spawn_rate” as “1”. This is why obtaining redstone is more difficult than obtaining diamond, as described in Section D.2.2.
Appendix D Task Details and Experiment Results
@@
D.1 Task Details
D.1.1 Context-Dependent Tasks
@@
D.1.2 Process-Dependent Tasks
@@
D.2 Success Rates of All Tasks
D.2.1 Context-Dependent Tasks
@@
Table 14: Detailed Performance on Context-Dependent Tasks. MethodA means the method uses the Active Perception strategy, and MethodG means the method uses the Fine-Grained Global Perception strategy. The parts with a gray background in the table represent the average success rate for the current level.
@@
D.2.2 Process-Dependent Tasks
@@
D.3 Ablation Study
D.3.1 Context-Dependent Tasks
We conduct ablation studies on the multi-modal large language model (MLLM) part within Context-Dependent Tasks in 15, comparing the performance outcomes of different MLLMs and different pre-trained visual encoders in the percipient.
D.3.2 Process-Dependent Tasks
@@
Appendix E Different Strategy of Active Perception
@@
Appendix F Applications
F.1 Obtain Diamond Pickaxe
@@ F.2 Discovery
F.3 Open-Ended Tasks